OpenStudy (anonymous):

Consider the function y = 3x5 – 25x3 + 60x + 1. Use the first derivative test to decide whether this function has a maximum at x = 1

OpenStudy (triciaal):

dy/dx = 15x^4 -75x^2 + 60 d^2y/dx = 60x^3 - 150x when x = 1 60 -150 = negative

OpenStudy (anonymous):

So is the derivative negative to the left of x = 1 and positive to the right of x = 1?

OpenStudy (anonymous):

Help?